. Titan Mining Corporation has 9 million shares of common stock outstanding and 200,000 bonds outstanding with 8 percent coupon rate, semiannual payments and par value $1,000 each. The common stock currently sells for $30 per share and has a beta of 1.20, and the bonds have 20 years to maturity and sell for 115 percent of par. The market risk premium is 6 percent, T-bills are 2 percent, and the company’s tax rate is 20 percent. If the company is evaluating a new investment project that has higher risk (beta of the project is 1.5) than the firm’s typical project, what rate should the firm use to discount the project’s cash flows?
Answer #1
1] Before tax cost of debt = YTM.
YTM using a financial calculator = 6.63%
After tax cost of debt = 6.63%*(1-20%) = 5.30%
2] Debt/Equity ratio:
Equity = 9000000*30 = $     27,00,00,000 54.00%
Debt = 200000*1150 = $     23,00,00,000 46.00%
Total capital $     50,00,00,000 100.00%
Debt/equity ratio = 46/54 = 0.8519
3] Levered beta for the project = 1.5*(1+0.8*0.8519) = 2.52
4] Cost of equity for the project per CAPM = 2%+2.52*6% = 17.12%
5] Discount rate for the project = 5.30%*46%+17.12%*54% = 11.68%